BRICS Notes Series, 2001

July 8, 2003

This document is also available as PostScript and DVI.

References

[NS-01-8] Abstract, PostScript, PDF.

Anders Mller and Michael I. Schwartzbach. The XML Revolution (Revised). December 2001. 186 pp. This revised and extended report superseeds the earlier BRICS Report NS-00-8.

[NS-01-7] Abstract, PDF.

Patrick Cousot, Lisbeth Fajstrup, Eric Goubault, Jeremy Gunawardena, Maurice Herlihy, Martin Rauen, and Vladimiro Sassone, editors. Preliminary Proceedings of the Workshop on Geometry and Topology in Concurrency Theory, GETCO ’01, (Aalborg, Denmark, August 25, 2001), August 2001. vi+97 pp.

[NS-01-6] Abstract, PDF.

Luca Aceto and Prakash Panangaden, editors. Preliminary Proceedings of the 8th International Workshop on Expressiveness in Concurrency, EXPRESS ’01, (Aalborg, Denmark, August 20, 2001), August 2001. vi+139 pp.

[NS-01-5] Abstract, PDF.

Flavio Corradini and Walter Vogler, editors. Preliminary Proceedings of the 2nd International Workshop on Models for Time-Critical Systems, MTCS ’01, (Aalborg, Denmark, August 25, 2001), August 2001. vi+ 127pp.

[NS-01-4] Abstract, PDF.

Ed Brinksma and Jan Tretmans, editors. Proceedings of the Workshop on Formal Approaches to Testing of Software, FATES ’01, (Aalborg, Denmark, August 25, 2001), August 2001. viii+156 pp.

[NS-01-3] Abstract.

Martin Hofmann, editor. Proceedings of the 3rd International Workshop on Implicit Computational Complexity, ICC ’01, (Aarhus, Denmark, May 20–21, 2001), May 2001. vi+144 pp.

[NS-01-2] Abstract, PostScript, PDF.

Stephen Brookes and Michael Mislove, editors. Preliminary Proceedings of the 17th Annual Conference on Mathematical Foundations of Programming Semantics, MFPS ’01, (Aarhus, Denmark, May 24–27, 2001), May 2001. viii+279 pp.

[NS-01-1] Abstract, PostScript, PDF. Nils Klarlund and Anders Mller. MONA Version 1.4 — User Manual. January 2001. 83 pp.